1. A soccer ball is kicked with a velocity of 29 m/s at  65° above the horizontal and lands on the roof of a 14.0 m building. 

What is the maximum height of the soccer ball?

If air resistance is ignored, what are the vertical velocity and resultant velocity when the ball hits the ground?



1
Expert's answer
2021-05-04T12:08:15-0400

"h_{max}=\frac{v_{0y}^2}{2g}=\frac{v^2\cdot\sin^265°}{2g}=\frac{29^2\cdot\sin^265°}{2\cdot9.81}=35.2\ (m/s)"


"v_{ver}=v_{0y}=v\cdot \sin65°=26.3\ (m/s)"


"v_{res}=v=29 \ (m/s)"